# Chinese Academy of Sciences, Beijing, China

### v2

@@ 5.1 Construction of two-to-one functions
The domain of a partial computable $g$ which is defined almost everywhere is a $\Pi^{0}_{2}$ class of measure 1 and includes all weakly 1-randoms. In general we only have $g\not\geq_{T}\emptyset^{\prime}$ so for some $r\not\geq_{T}\emptyset^{\prime}$ we use weakly $r$-randoms: reals that are members of every $\Sigma^{0}_{1}(r)$ class of measure 1.
@@
Let $r\in 2^{\omega}$. A real $w$ is called $r$-generic if for every $\Sigma^{0}_{1}(r)$ class $G\subseteq 2^{\omega}$
@@
By [11] and [12] ([4, Theorem 8.11.7]) respectively, for each $r$:
@@
Given $r\geq_{T}g$, we adapt §5.2 by choosing $y_{i},z$ so that the $y_{i}\oplus z$ are weakly $r$-random. We construct them from the $y,w$ given by:
@@
Let $w$ be $r$-generic and $y$ be $(w\oplus r)^{\prime}$-random. Then
  • $w\oplus r\not\geq_{T}\emptyset^{\prime}$
  • $w$ is weakly $r$-random
  • $y$ is weakly 2-random relative to $w\oplus r$.
By the latter two clauses and [1, Corollary 2.1] we get (i).
By [20, Proposition 2.2] each $w^{n}$ is $r$-generic. Since $r$-generics are not randoms we get (ii). Since $y$ is weakly 2-random relative to $w\oplus r$, by [12] ([4, Theorem 7.2.7]), $y$ is not a member of any $\Pi^{0}_{2}(w\oplus r)$ null class. The latter also holds for $\Sigma^{0}_{3}(w\oplus r)$ null classes which are countable unions of $\Pi^{0}_{2}(w\oplus r)$ classes. Since $w\oplus r\not\geq_{T}\emptyset^{\prime}$, by [4, Corollary 8.12.2] the class
@@
has measure 0. Since $G$ is a $\Sigma^{0}_{3}(w\oplus r)$ class we get $y\not\in G$ so (iii) holds. ∎
Fix $w,y$ as in Lemma 5.8 and let $U$ be a member of a universal Martin-Löf test with effective enumeration $(U_{s})$, $y\not\in U$ and $\forall n,\ w^{n}\in U$. For each $n$ we can $(w\oplus y)$-effectively define a weakly $r$-random $z=z_{n}$ consisting of the columns of $w$ except for the $n$th column which is $y$. This suggests defining $E^{z}_{s}(i)$ in the template of §5.1 in terms of the memberships $z^{i}\in U$.
@@
so $k^{z}_{s_{n}}$ does not get $z$-permission. In §5.2 permissions of $k_{s}^{z}$ at $s$ depended entirely on the value of $k_{s}^{z}$ so (9) was achieved by defining $s_{n},z$ with $k_{s_{n}}^{z}=n$. This is no longer possible as we do not have control over the stages where the $w^{i}$ appear in $U$. The solution is to define permissions in terms of
@@
Construction and verification
Given $z$ let $k^{z}_{s}$ be the non-decreasing counter with $k^{z}_{0}=0$ and
@@
enumerates $\mathbb{N}$, omitting $k^{z}_{s}$ as long as it is not updated and including it if an update occurs. Updates of $k^{z}$ coincide with updates of $d^{z}$ and are due to:
  • $d^{z}_{s}\in\emptyset^{\prime}_{s}$ which we call $\emptyset^{\prime}$-permission of $k^{z}_{s}$ at $s+1$; or
@@
Both $d^{z},k^{z}$ are non-decreasing and $d^{z}$ increases by at most 1. The reason that $k^{z}$ is allowed to skip numbers is so that the range of $p^{z}$ misses at most one number $m$, which happens exactly when $\lim_{s}k^{z}_{s}=m$. So
$k^{z}_{s}$ becomes some $p^{z}_{t}$ iff $d^{z}_{s}$ receives permission at some stage.
@@
Given $r\not\geq_{T}\emptyset^{\prime}$ let $y,w$ be as in Lemma 5.8. Effectively in $y\oplus w,n$ we can define $z$ and $s$ with
@@
    $d^{z}_{s}=n\ \wedge\ \ \big{(}{\lim_{t}k^{z}_{t}<\infty\iff\lim_{t}k^{z}_{t}=k^{z}_{s}\iff n\not\in\emptyset^{\prime}}\big{)}$
@@
so $h^{z}(x)$ outputs the bits of $x$ in some order determined by $(\emptyset^{\prime}_{s})$, $z$ with the exception of $\lim_{s}k^{z}_{s}$ in the case that the limit is finite. Then (i), (ii) of §5.1 hold and $f$ is two-to-one. Since $s\mapsto p^{z}_{s}$ is injective, Lemma 5.5 shows that $f$ is a computable random-preserving surjection.
Assuming that $g:\subseteq 2^{\omega}\to 2^{\omega}$ inverts $f$ with probability 1, we will show that $g\geq_{T}\emptyset^{\prime}$.
@@
fix $y_{0},v,w$ as in Lemma 5.8 for $y=y_{0}$ and $r$
so $y_{0}\oplus w\oplus r\not\geq_{T}\emptyset^{\prime}$.
To decide if $n\in\emptyset^{\prime}$, from $y_{0}\oplus w\oplus r$ we effectively compute $z,s$ as in Lemma 5.9. Let $k:=k^{z}_{s}$ be the potential finite limit of $k^{z}$. Then $y_{0}\oplus z$ is weakly $r$-random, $d^{z}_{s}=n$ and
$\forall t>s\ \big{(}{n\in\emptyset^{\prime}_{t}\ \Rightarrow\ p^{z}_{t}=k}\big{)}.$
So by the definition of $f$, for each $x,y$:
$\big{(}{f(x\oplus z)=y\oplus z\ \wedge\ \ t>s\ \wedge\ \ n\in\emptyset^{\prime}_{t}}\big{)}\ \Rightarrow\ y(t)=x(k).$ (10)
Now we show $n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}$ where
$g(y_{0}\oplus z;2k)\downarrow\hskip 8.5359pt\textrm{with oracle-use $u$.}$ (11)
Since $y_{0}\oplus z$ is weakly $r$-random, by Lemma 5.6 and we have $g(y_{0}\oplus z)\downarrow$ so $u$ is well-defined. For a contradiction let $t$ be such that
$t>\max\{u,s\}\ \wedge\ \ n\in\emptyset^{\prime}_{t}-\emptyset^{\prime}_{t-1}$
and let $y_{1}$ be the real with $y_{1}(i)=y_{0}(i)\iff i\neq t$. Then $y_{1}\oplus z$ is weakly $r$-random, so by Lemma 5.6, for $i=0,1$
$g(y_{i}\oplus z)\downarrow\ \wedge\ \ f(g(y_{i}\oplus z))=y_{i}\oplus z.$ (12)
Since $t>u$ we have $(y_{0}\oplus z)\upharpoonright_{u}\prec y_{1}\oplus z$ so by (11):
@@
which contradicts (13) since $y_{0}(t)\neq y_{1}(t)$. We have shown that
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}$
which gives a decision procedure for $\emptyset^{\prime}$ from $y_{0}\oplus w\oplus r$, contradicting the choice of $y_{0},w,r$ such that $y_{0}\oplus w\oplus r\not\geq_{T}\emptyset^{\prime}$. We conclude that $g\geq_{T}\emptyset^{\prime}$ so every almost everywhere inversion of $f$ computes $\emptyset^{\prime}$. The same argument applies in the case that $g$ inverts $f$ almost everywhere in a neighborhood $\llbracket\sigma\rrbracket$.

### v3

@@ 4.2 Almost everywhere probabilistic inversions
The domain of a partial computable $g$ which is defined almost everywhere is a $\Pi^{0}_{2}$ class of measure 1 and includes all weakly randoms. In general we only have $g\not\geq_{T}\emptyset^{\prime}$ so for some $r\not\geq_{T}\emptyset^{\prime}$ we use weakly $r$-randoms: reals that are members of every $\Sigma^{0}_{1}(r)$ class of measure 1.
@@
At this point we need some facts about generic reals.
Genericity is a topological form of typicality which has been extensively studied along with randomness in computability [17, 4] and computational complexity [29, 28, 1, 23]. Generic reals avoid every definable meager (as in Baire category) set, as algorithmically random reals avoid definable null set. The level of definability determines the strength of genericity or randomness.
Definition 4.5 (Jockusch [17]).
Given $r,w\in 2^{\omega}$, if
@@
for every $\Sigma^{0}_{1}(r)$ class $G\subseteq 2^{\omega}$ we say that $w$ is $r$-generic.
By [17] and [18] (see [8, Theorem 8.11.7]) respectively, for each $r$:
  1.    (a)
@@
By the analogue of (6) for genericity [30, Proposition 2.2] we have
if $w$ is $r$-generic then $w^{n}$ is $r$-generic for each $n$. (11)
By [3, Corollary 2.1] one implication of (6) holds for weak randomness:
if $w$ is weakly $r$-random and $y$ is weakly $w\oplus r$-random then $y\oplus w$ is weakly $r$-random. (12)
Recall that incomputable sets cannot be computed probabilistically with non-zero probability [7]. A relativization of this fact is
$y\not\geq_{T}r\implies\mu\big{(}{\{\hskip 0.85355pt{x}:{x\oplus y\geq_{T}r}\hskip 0.85355pt\}}\big{)}=0$ (13)
which can be found in [8, Corollary 8.12.2].
Given $r\geq_{T}g$, we adapt the argument of §4.1 by choosing $y_{i},z$ so that the $y_{i}\oplus z$ are weakly $r$-random. We construct them from the $y,w$ given by:
@@
Let $w$ be $r$-generic and $y$ be $(w\oplus r)^{\prime}$-random so
$y$ is weakly 2-random relative to $w\oplus r$ (14)
and by (b), $w$ is weakly $r$-random. This, combined with (12), gives (i).
By (11) each $w^{n}$ is $r$-generic and by (b) not random, hence (ii).
By (a) we have $w\oplus r\not\geq_{T}\emptyset^{\prime}$ so by (13) the $\Sigma^{0}_{3}(w\oplus r)$ class
@@
is null. By (14) and (2) we get $y\not\in G$ and (iii) holds. ∎
Fix $w,y$ as in Lemma 4.7 and let $(U_{s})$ be an effective enumeration of a member $U$ of a universal Martin-Löf test with $y\not\in U$ and $\forall n,\ w^{n}\in U$.
For each $n$ we $(y\oplus w)$-effectively define a real $z$ consisting of the columns of $w$ except for the $n$th column which is $y$. Then $y\oplus w$ is weakly $r$-random by (12). Since $z$ copies $y\oplus w$ from a computable array of indices, $z$ is also weakly $r$-random. This suggests defining $E^{z}_{s}(i)$ in the template of §4.1 in terms of the memberships $z^{i}\in U$ of the $i$th column $z^{i}$ of $z$.
@@
so $k^{z}_{s_{n}}$ does not get $z$-permission. In §4.1 permissions of $k_{s}^{z}$ at $s$ depended entirely on the value of $k_{s}^{z}$ so (15) was achieved by defining $s_{n},z$ with $k_{s_{n}}^{z}=n$. This is no longer possible as we do not have control over the stages $s$ where the $w^{i}$ appear in $U$. The solution is to define permissions in terms of
@@
Parameters. Given $z$ let $k^{z}_{s}$ be the non-decreasing counter with $k^{z}_{0}=0$ and
@@
enumerates $\mathbb{N}$, omitting $k^{z}_{s}$ as long as it is not updated and including it otherwise. Updates of $k^{z}$ coincide with those of $d^{z}$ and are due to one of:
  • $d^{z}_{s}\in\emptyset^{\prime}_{s}$ which we call $\emptyset^{\prime}$-permission of $k^{z}_{s}$ at $s+1$
@@
Both $d^{z},k^{z}$ are non-decreasing and $d^{z}$ increases by at most 1. The reason that $k^{z}$ is allowed to skip numbers is so that the range of $p^{z}$ misses at most one number $m$, which happens exactly when $\lim_{s}k^{z}_{s}=m$.
So $k^{z}_{s}$ becomes some $p^{z}_{t}$ iff $d^{z}_{s}$ receives permission at a stage $t>s$.
@@
Given $r\not\geq_{T}\emptyset^{\prime}$ let $y,w$ be as in Lemma 4.7. Effectively in $y\oplus w$ and $n$ we can define $z$ and $s$ such that
@@
    $d^{z}_{s}=n$ and $\big{(}{\lim_{t}k^{z}_{t}<\infty\iff\lim_{t}k^{z}_{t}=k^{z}_{s}\iff n\not\in\emptyset^{\prime}}\big{)}$
@@
We are now ready to prove:
Theorem 4.9.
There is a total computable $f:2^{\omega}\to 2^{\omega}$ such that:
  • $f$ is a two-to-one random-preserving surjection
  • for each partial $g\not\geq_{T}\emptyset^{\prime}$, with positive probability, $g$ fails to invert $f$
and the latter holds for the restriction of $f$ in any cylinder $\llbracket\sigma\rrbracket$.
Proof..
@@
so $h^{z}(x)$ outputs the bits of $x$ in some order determined by $(\emptyset^{\prime}_{s})$, $z$ with the exception of $\lim_{s}k^{z}_{s}$ when this limit is finite. Then (i), (ii) of §4.1 hold and $f$ is two-to-one. Since $s\mapsto p^{z}_{s}$ is injective, Lemma 4.2 shows that $f$ is a computable random-preserving surjection.
Assuming $g:\subseteq 2^{\omega}\to 2^{\omega}$ inverts $f$ with probability 1, we show that $g\geq_{T}\emptyset^{\prime}$.
@@
fix $y_{0},v,w$ as in Lemma 4.7 for $y=y_{0}$, so $y_{0}\oplus w\oplus r\not\geq_{T}\emptyset^{\prime}$.
Fix $n$. To decide if $n\in\emptyset^{\prime}$, from $y_{0}\oplus w\oplus r$ we
  • effectively compute $z,s$ as in Lemma 4.8
  • let $k:=k^{z}_{s}$ be the potential finite limit of $k^{z}$
so $y_{0}\oplus z$ is weakly $r$-random, $d^{z}_{s}=n$ and $\forall t>s\ \big{(}{n\in\emptyset^{\prime}_{t}\implies p^{z}_{t}=k}\big{)}$.
By the definition of $f$, for each $x,y$:
$\big{(}{f(x\oplus z)=y\oplus z\ \wedge\ \ t>s\ \wedge\ \ n\in\emptyset^{\prime}_{t}}\big{)}\implies y(t)=x(k).$ (16)
Since $y_{0}\oplus z$ is weakly $r$-random, by Lemma 4.4 we have $g(y_{0}\oplus z)\downarrow$.
Letting $u$ be the oracle-use of $g(y_{0}\oplus z;2k)\downarrow$ we claim that
$n\in\emptyset^{\prime}\iff n\in\emptyset^{\prime}_{\max\{u,s\}}.$ (17)
Otherwise there is $t>\max\{u,s\}$ with $n\in\emptyset^{\prime}_{t}-\emptyset^{\prime}_{t-1}$. Let $y_{1}$ be the real with
$y_{1}(i)=y_{0}(i)\iff i\neq t.$
Then $y_{1}\oplus z$ is weakly $r$-random, so by Lemma 4.4, for $i=0,1$
$g(y_{i}\oplus z)\downarrow\hskip 8.5359pt\textrm{and}\hskip 11.38092ptf(g(y_{i}\oplus z))=y_{i}\oplus z.$ (18)
Since $t>u$ we have $(y_{0}\oplus z)\upharpoonright_{u}\prec y_{1}\oplus z$ so by (17) we get
@@
which contradict (19) since $y_{0}(t)\neq y_{1}(t)$. This concludes the proof of (17).
By (17) we get $\emptyset^{\prime}\leq_{T}y_{0}\oplus w\oplus r$ which contradicts the hypothesis that $y_{0}\oplus w\oplus r\not\geq_{T}\emptyset^{\prime}$. We conclude that $g\geq_{T}\emptyset^{\prime}$ so every almost everywhere inversion of $f$ computes $\emptyset^{\prime}$. The same argument applies in the case that $g$ inverts $f$ almost everywhere in a cylinder $\llbracket\sigma\rrbracket$. ∎
5 Conclusion
We constructed a one-way function on the reals, which is the analogue of the one-way functions in computational complexity formulated by Levin in [20]. We argued that Levin’s definition is the correct analogue of the one-way functions from computational complexity, over weaker alternatives. An analysis of alternatives was conducted in [5], along with a study of the oracles needed to probabilistically invert one-way real functions.
Our result was based on a general framework for constructing permutations of selected bits of the input, which was adapted toward understanding the extent to which one-way functions can be injective. Applications often require additional ideas, but can yield analogues of significant properties in computational complexity, such as collision-resistance [2].
Despite the non-trivial adaptations that are often required, all currently known one-way real functions (as defined in [20]) are permutations of selected bits of the input and can thus be viewed as applications of our framework. The question therefore arises with respect to the generality of this framework. This is relevant to the existence of a partial computable one-way injection, which is currently unknown.